# Statistical Guarantees Of False Discovery Rate In Medical Instance Segmentation Tasks Based on Conformal Risk Control

### v1

@@ IV-D Evaluation Metrics
IV-D1 False Discovery Rate (FDR)
@@
IV-D2 False Negative Rate (FNR)
@@
IV-D3 Conformal Prediction Metrics
@@ IV-F Practical Steps
In the experiment, we select FDR as the core metric for the calibration loss function, focusing on the precision of segmentation results. This choice ensures that the output of the prediction model remains as accurate as possible, avoiding excessive false positives (false discoveries), even if it might miss some target regions. We use both Mask R-CNN [1] and BlendMask [2] models and generate 10 test results through cross-validation.
@@
In the experimental design, we focus on FNR as the primary goal of the calibration loss function, aiming to maximize the coverage of segmentation results, i.e., covering as many correct tumor regions as possible. This choice ensures that the prediction model captures all possible target regions comprehensively, even at the cost of increasing some false positives (false negatives). To verify the effectiveness of this method, we use both Mask R-CNN [1] and BlendMask [2] models and generate 10 test results through cross-validation.

### v2

@@ 4.4 Evaluation Metrics
4.4.1 False Discovery Rate (FDR)
@@
4.4.2 False Negative Rate (FNR)
@@
4.4.3 Conformal Prediction Metrics
@@ 4.6 Practical Steps
In the experiment, we select FDR as the core metric for the calibration loss function, focusing on the precision of segmentation results. This choice ensures that the output of the prediction model remains as accurate as possible, avoiding excessive false positives (false discoveries), even if it might miss some target regions. We use both Mask R-CNN He et al. (2017) and BlendMask Chen et al. (2020) models and generate 10 test results through cross-validation.
@@
In the experimental design, we focus on FNR as the primary goal of the calibration loss function, aiming to maximize the coverage of segmentation results, i.e., covering as many correct tumor regions as possible. This choice ensures that the prediction model captures all possible target regions comprehensively, even at the cost of increasing some false positives (false negatives). To verify the effectiveness of this method, we use both Mask R-CNN He et al. (2017) and BlendMask Chen et al. (2020) models and generate 10 test results through cross-validation.